• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Humensky

    Log entry time 16:08:49 on November 1,1998

    Entry number 4407

    This entry is a followup to: 4399

    keyword=beam charge asymmetry

    I made a typo in my prior note, 4399... The monitor to follow is Q10, not Q1. Here's the complete text again, corrected:

    The plots below are time histories of the beam current asymmetry since ~2:30am 11/1/98. In order from top to bottom, the asymmetries are measured by Q1, Q10, Clock. We want to see Q10<100ppm.

    Q10 has mostly been at 60-70ppm, except for a couple excursions to 150ppm. Run 2262 was very short, so its large asymmetry is dominated by the timing error.

    This should be checked each run, and if it is seen moving upwards of 100ppm, Matt Poelker or Etienne Burtin should be called about trying to correct it.